# From Commands to Prompts: LLM-based Semantic File System for AIOS

### v1

@@ 4.4 LSFS Parser
4.5 The Interaction between Modules
In Figure 4, we present the examples to demonstrate how components of LSFS interact with each other to achieve different functionalities. The upper section of Figure 4 depicts the workflow of the retrieve-summary API, while the lower section outlines the workflows of the change-summary API and rollback API. In the upper part of Figure 4, the LSFS parser decodes prompts into API calls with API name and API arguments. Then LSFS executes the API to check vector database to get results. This API also provides user-interaction interface for the users to verify results. After the verification, the content will be summarized by leveraging LLM. In the lower part of Figure 4, when a modification request is submitted, the LSFS parser decodes the file information (name and location) that is to be modified. The LSFS then modifies the semantic changes in both the vector database and the files stored in the disk. Meanwhile, the supervisor of the LSFS is kept running to ensure consistency between the semantic index of files in the LSFS and the files stored in the disk. Upon updated, the summarization API compares the file contents before and after the change to generate a detailed change log. Additionally, the API stores the pre-modification content in the version recorder. If a rollback is requested, the API retrieves the specific version from the version recorder and synchronizes it in both the LSFS and files in the disk.
@@ 5.1 RQ1: Effectiveness of LSFS Parser
For RQ1, we assess the accuracy of the LSFS parser in translating user natural language prompt into executable LSFS API calls. We evaluate the accuracy of LSFS parser with 30 different samples for each API on different LLM backbones, i.e., Gemmi-1.5-Flash, GPT-4o-mini, Qwen-2, and Gemma-2. The results, illustrated in Figure 5, reveal that the LSFS parser performs exceptionally well on parsing prompts related to change-summary API and link API (for which the semantic information in the user prompt is relatively simple), achieving high accuracy across all LLMs, with some reaching 100%
@@
accuracy. For more complex prompts, such as those intended for the rollback API and retrieve-summary API (for which the semantic information in the user prompt is complex), accuracy remains above 85% for most models, except for Gemma-2. The average parsing accuracy reaches 90%. These results show that the LSFS parser can effectively translate natural language information into executable API calls, showcasing its reliability in diverse scenarios. In all cases, the translated command is provided to users for approval before real execution, ensuring safety when executing commands in the file system.
@@ 5.2 RQ2: Analysis of LSFS in Semantic File Management Tasks
In our experiments, we compare the performance of using LSFS and without using LSFS under the same LLM backbone. The details of the prompts we use for the comparison are in the Appendix  B. Specifically, we use Gemini-1.5-flash and GPT-4o-mini as the LLM backbone, respectively, for the comparison. As shown in Table 2, using LSFS to implement the retrieval function significantly enhances both the accuracy and the efficiency compared to only leveraging LLM for retrieval without using LSFS. As file number increases, the retrieval accuracy tends to drop significantly when using LLM for retrieval without LSFS. This is because more files can lead to longer context for the LLM, which degrades the LLM’s performance of identifying information in the long context. By contrast, using LSFS can still achieve good retrieval accuracy and have much better retrieval efficiency when file number increases, because lsfs replaces the reasoning process of LLM by using keyword matching and semantic similarity matching, it saves a lot of time and avoids the errors of LLM when facing complex input text.
Figure 6: A file summarization agent instance to illustrate the comparison between LLM-leveraged methods without using LSFS and using LSFS.
[…]
@@
LSFS supports semantic file rollback, which enables the restoration of a file to a particular version specified by the time requested by the user or number of versions, recorded by the Version Recorder. We vary the the number of rollback versions and calculate their corresponding rollback time, to evaluate the stability and efficiency of the version rollback process, The results, shown in Figure 7, illustrate the consistency in the time consumed during version rollbacks. We use Gemmi-1.5-Flash, GPT-4o-mini, Qwen2 as the LLM backbones for the experiments. In our experiment, we set the file version rollback process to begin from the newest one version, with a rollback occurring every last 5 versions. The experiment continues until a maximum of 40 versions had been rolled back. Each rollback is simultaneously updated in both the LSFS and file of the computer’s local disk. As shown in the Figure 7, across all three LLM backbones, the rollback time does not increase exponentially with the number of versions rolled back. Instead, it tends to plateau. This is because the rollback API stores each version of the file independently, thereby avoiding long rollback paths to implement rollbacks and achieving a stable rollback time.
@@ 5.3 RQ3: Analysis of LSFS in Non-Semantic File Management Tasks
In this section, we compare LSFS and traditional file system in keyword-based file retrieval task. The task is to use keywords existing in the filename or file content to retrieve files. We build a hierarchical file folder with file numbers as 10, 20, and 40, respectively, for this task. We use two types of retrieval prompts, i.e., single-condition and multi-condition, to evaluate LSFS and traditional file systems to retrieve relevant files containing specific keywords. The details of how we construct prompts for this task can be seen at Appendix B. We consider the following methods as the retrieval baselines in the traditional file system. It is important to note that the original grep command can only deal with plain text files, such as .txt and .md files, and cannot support binary files, such as .pdf and .doc files. Therefore, we make two enhanced versions, named as TFS-grep and TFS-grep* to make the comparison.
  • TFS search window: We use the default search window in the computer’s file folder to retrieve files (i.e., Spotlight in MacOS) which supports retrieving by keywords in the file content.
Table 3: Comparison between LSFS and methods in the traditional file system (TFS) in retrieving files by keywords in the names and content of files.
@@
We use precision, recall and F1-score to measure the retrieval performance. From the results presented in Table 3, we can find that LSFS outperforms TFS search window and TFS-grep, only second to the TFS-grep*. We find that the built-in retrieval tool in the TFS (e.g., the system search window) can not generate stable retrieval results, although it has a higher recall. Due to the fuzzy search feature in the built-in search window, it can easily retrieve inaccurate results for which the retrieved file content can only match part of the keywords. For example, if we search for an article written by John Smith, any other article with a name of John can be returned as a search result, thus the results may often include many irrelevant results, which complicates the process for users to filter through them. The TFS-grep* command, although achieving perfect results, still has several limitations. First, the commands can be too difficult to construct, especially when the file retrieval queries have multiple conditions. For instance, when a user requests to retrieve all files containing the keywords A and B, the command would be as follows: find /path -type f -exec grep -l "keyword1" $\\\backslash\;$ -exec grep -l "keyword2" $\backslash\;$. Second, since the grep command itself does not support retrieval of binary files, it is necessary to manually adjust the format of each file, which is time-consuming and greatly reduces the efficiency of the retrieval process for users. Our LSFS can retrieve all types of text files, from plain text to binary files, while maintaining high precision and recall. The LSFS read operation is capable of processing both plain text and binary text, converting them into the system’s vector database. This enables seamless retrieval operations across various types of files. Furthermore, LSFS allows users to describe their retrieval tasks in natural language, eliminating the need to write complex commands.
@@
For the file sharing task, we evaluate whether a system can output a shareable link with an expiration time according to the prompts. Specifically, we compare LSFS with four different baselines as following.
  • GPT-4o-mini: We employ GPT-4o-mini as the LLM backbone to generate code for creating links of the target file, followed by using the Python compiler to verify the code’s validity.
@@
Method Success rate of generating sharable links (#20) Code Generation Rate Link Generation Rate Link Validness Rate Final Success Rate Gemini-1.5-flash 65% 45% 45% 10% GPT-4o-mini 60% 35% 30% 5% AutoGPT 50% 45% 15% 5% Code Interpreter 100% 75% 65% 0% LSFS 100% 100% 100% 100%
@@ 6 Conclusions
In this paper, we present an LLM-based semantic file system (LSFS), which offers advancement over traditional file systems by enabling files to be stored and managed based on their semantic information. This enhancement improves the system’s ability to comprehend and utilize the semantics embedded in file contents. Additionally, we introduce a series of reusable semantic syscalls and a framework for mapping natural language into LSFS parameters. These innovations provide a foundation for future research and development in the area of semantic file management systems. Looking ahead, future work can further explore and expand the application of LSFS across a wider scope of user environments and integrate semantic file management into everyday computing. This can help to simplify user operations and pave the way for a more intelligent and user-friendly operating system.
7 Acknowledgement
We thank Minghua Ma and Sam Lin for the useful discussions with them.
@@ Appendix A The Implementation Details of Syscall
keywords_retrieve() This syscall implements LSFS’s keyword search functionality, retrieving all files within a specified directory that contain a given keyword. The arguments passed include the LSFS path, directory name, keyword, and matching condition. The LSFS path and keyword are Positional Parameters, while the directory name and condition are Default Parameters. If a directory name is provided, the syscall retrieves files within that directory that match the keyword; otherwise, it searches across the entire system. To match multiple keywords, the matching condition must be passed, specifying the relationship between keywords, such as and or or. The search results return a list of file names and a list of file contents.
@@
file_join() This syscall facilitates the connection of two files, with the parameters including the LSFS path, the directory name and filename of both files, and the connection conditions. The Default Parameters are the destination directory for file 2 and the connection condition, while the LSFS path, directory names, and filenames are Positional Parameters. If the destination directory for file 2 does not exist, the two files will be placed in the same directory. If the destination directory exists, the files will be placed in their respective directories. If the join condition is set to new, the syscall will preserve the original files, concatenate their text contents, and create a new file in the destination directory of file 1. The new file’s name will be a combination of the two target filenames. If the join condition is not new, the contents of file 2 will be appended directly to the contents of file 1, and file 2 will then be deleted.
Appendix B Details of keyword-based and semantic retrieval.
Task Task Example Method Instruction
Keyword-based Retrieval (Single-condition) Find papers in the computer-vision category authored by Emily Zhang. LLM w/o LSFS At current step, you need to judge if the input paper satisfy [retrieve condition]. If yes, you should summarize the paper, if no you don’t need to output anything. The paper is [file content]
Keyword-based Retrieval (Multi-condition) Find papers from either Cambridge University or Columbia University. LLM w/o LSFS At current step, you need to judge if the input paper satisfy [retrieve condition]. If yes, you should summarize the paper, if no you don’t need to output anything. The paper is [file content]
@@
Appendix C The Instruction of API
Table 7: Some examples of instruction of API in Section 4.3. For every API, we provide different instructions in different task condition. The instruction of retrieve-summary API is in Appendix B.
API Type Instruction
w/o directory: Modify /xxxx/xxxx.txt to contain change-file.
w/o period of validity: Generate a link for system-architecture.
Appendix D Experiment Details of File Sharing
In this module, we show some examples that the LLM execution results in a failure to correctly generate shareable links. We give each LLM the following input:You are good at writing code, please write code to generate shared links for the file ’path’.
D.1 The code cannot generate link
In the experiment, the code generated by LLMs may not produce the correct link or link address. For instance, even after successfully installing the required file packages, the following code block demonstrates that the generated link does not direct to the intended target file. The pseudo-code is in Algorithm 1
Algorithm 1 Pseudo-code of D.1.
@@ D.1 The code cannot generate link
D.2 The code can only generate local link
@@
Algorithm 2 Pseudo-code of D.2.
@@
D.3 The code can generate shareable link
@@
Appendix E Further Discussions
@@
The core functionality of LSFS lies in its keyword and semantic retrieval capabilities. In future work, these retrieval methods can be further optimized by integrating more advanced and precise algorithms, enhancing the overall accuracy and effectiveness of LSFS’s retrieval performance.

### v5

@@ 5.1 RQ1: Effectiveness of LSFS Parser
For RQ1, we assess the accuracy of the LSFS parser in translating user natural language prompt into executable LSFS API calls. We evaluate the accuracy of LSFS parser with 30 different samples for each API on different LLM backbones, i.e., Gemmi-1.5-Flash, GPT-4o-mini, Qwen-2, and Gemma-2. The results, illustrated in Figure 5, reveal that the LSFS parser performs exceptionally well on parsing prompts related to change-summary API and link API (for which the semantic information in the user prompt is relatively simple), achieving higher accuracy across all LLMs (i.e., over 90%), where GPT-4o-mini and Qwen-2 all reach 100% accuracy. For more complex prompts, such as those intended for the rollback API and retrieve-summary API (for which the semantic information in the user prompt is complex), accuracy remains above 85% for most models, except for Gemma-2.
@@
The average parsing accuracy reaches 90%. These results show that the LSFS parser can effectively parse natural language information into executable API calls, showcasing its reliability in diverse scenarios. For safety consideration, in all cases, the parsed API calls are provided to users for confirmation and approval before execution, avoiding irreversible file operations like deleting files or directories.
@@ 5.2 RQ2: Analysis of LSFS in Semantic File Management Tasks
To answer RQ2, we evaluate the performance LSFS on semantic file management tasks (i.e., semantic file retrieval and semantic file rollback).
@@
In our experiments, we compare the performance of using LSFS and without using LSFS under the same LLM backbone. The details of the prompts we use for the comparison are in the Appendix D. Specifically, we use Gemini-1.5-flash and GPT-4o-mini as the LLM backbone, respectively, for the comparison. As shown in Table 2, using LSFS to implement the retrieval function significantly enhances both the accuracy and the efficiency compared to only leveraging LLM for retrieval without using LSFS. As file number increases, the retrieval accuracy tends to drop significantly when using LLM for retrieval without LSFS. This is because more files can lead to longer context for the LLM, which degrades the performance of LLM of identifying information in the long context. By contrast, using LSFS can still achieve good retrieval accuracy and have much better retrieval efficiency when file number increases, because lsfs replaces the reasoning process of LLM by using keyword matching and semantic similarity matching, it saves a lot of time and avoids the errors of LLM when facing complex input text. The concrete example of API with LSFS and API without LSFS are in Appendix F.
@@
LSFS supports semantic file rollback, which enables the restoration of a file to a particular version specified by the time requested by the user or number of versions, recorded by the Version Recorder. We vary the the number of rollback versions and calculate their corresponding rollback time, to evaluate the stability and efficiency of the version rollback process, The results, shown in Figure 6, illustrate the consistency in the time consumed during version rollbacks. We use Gemmi-1.5-Flash, GPT-4o-mini, Qwen2 as the LLM backbones for the experiments. In our experiment, we rollback file with versions the range from 5 to 40, using increments of 5. Each rollback is simultaneously updated in both the LSFS and file stored in the disk. As shown in the Figure 6, across all three LLM backbones, the rollback time does not increase exponentially with the number of versions rolled back. Instead, it tends to plateau with a stable rollback time <<< 1 min even if file number increases, suggesting the scalability of semantic file rollback supported by LSFS.
@@ 5.3 RQ3: Analysis of LSFS in Non-Semantic File Management Tasks
In this section, we compare LSFS and traditional file system in keyword-based file retrieval task. The task is to use keywords existing in the filename or file content to retrieve files. We build a hierarchical file folder with file numbers as 10, 20, and 40, respectively, for this task. We use two types of retrieval prompts, i.e., single-condition and multi-condition, to evaluate LSFS and traditional file systems to retrieve relevant files containing specific keywords. The details of how we construct prompts for this task can be seen at Appendix D. We consider the following methods as the retrieval baselines in the traditional file system. It is important to note that the original grep command can only deal with plain text files, such as .txt and .md files, and cannot support binary files, such as .pdf and .doc files. Therefore, we make two enhanced versions, named as TFS-grep and TFS-grep* to make the comparison. The experimental setup is presented in Appendix E.
Table 3: Comparison between LSFS and methods in the traditional file system (TFS) in retrieving files by keywords that match names and content of files.
@@
We use precision, recall and F1-score to measure the retrieval performance. From the results presented in Table 3, we can find that LSFS outperforms TFS search window and TFS-grep, only second to the TFS-grep*. We find that the built-in retrieval tool in the TFS (e.g., the system search window) can not generate stable retrieval results, although it has a higher recall. Due to the fuzzy search feature in the built-in search window, it can easily retrieve inaccurate results for which the retrieved file content can only match part of the keywords. For example, if we search for an article written by John Smith, any other article with a name of John can be returned as a search result, thus the results may often include many irrelevant results, which complicates the process for users to filter through them. The TFS-grep* command, although achieving perfect results, still has several limitations. First, the commands can be too difficult to construct, especially when the file retrieval queries have multiple conditions. For instance, when a user requests to retrieve all files containing the keywords A and B, the command would be as follows: find /path -type f -exec grep -l ”keyword1” $\\\backslash\;$ -exec grep -l ”keyword2” $\backslash\;$. Second, since the grep command itself does not support retrieval of binary files, it is necessary to manually adjust the format of each file, which is time-consuming and greatly reduces the efficiency of the retrieval process for users. Our LSFS can retrieve all types of text files, from plain text to binary files, while maintaining high precision and recall. The LSFS read operation is capable of processing both plain text and binary text, converting them into the vector database of system. This enables seamless retrieval operations across various types of files. Furthermore, LSFS allows users to describe their retrieval tasks in natural language, eliminating the need to write complex commands.
@@
For the file sharing task, we evaluate whether a system can output a shareable link with an expiration time according to the prompts. Specifically, we compare LSFS with four different baselines which details are in Appendix E.
@@
[…]
@@ 6 Conclusions
In this paper, we present an LLM-based semantic file system (LSFS), which offers advancement over traditional file systems by enabling files to be stored and managed based on their semantic information. This enhancement improves the ability of system to comprehend and utilize the semantics embedded in file contents. Additionally, we introduce a series of reusable semantic syscalls and a framework for mapping natural language into LSFS parameters. These innovations provide a foundation for future research and development in the area of semantic file management systems. Looking ahead, future work can further explore and expand the application of LSFS across a wider scope of user environments and integrate semantic file management into everyday computing. This can help to simplify user operations and pave the way for a more intelligent and user-friendly operating system.
@@
Roadmap: Section A introduces the detailed implementations of LSFS sycalls. Section B introduces the detailed implementations of LSFS APIs. Section C shows concrete prompt examples to execute different APIs. Section D shows the details of semantic file retrieval and keyword-based file retrieval. Section F shows the results of a case study of comparing methods using LSFS and without using LSFS in semantic file retrieval task. Section G presents analysis of file sharing tasks with pseudo-code examples of the file sharing process. Section H discusses future work that can be built on LSFS.
@@ Appendix A The Implementation Details of Syscall
keywords_retrieve() This syscall implements keyword search functionality of LSFS, retrieving all files within a specified directory that contain a given keyword. The arguments passed include the LSFS path, directory name, keyword, and matching condition. The LSFS path and keyword are Positional Parameters, while the directory name and condition are Default Parameters. If a directory name is provided, the syscall retrieves files within that directory that match the keyword; otherwise, it searches across the entire system. To match multiple keywords, the matching condition must be passed, specifying the relationship between keywords, such as and or or. The search results return a list of file names and a list of file contents.
@@
file_join() This syscall facilitates the connection of two files, with the parameters including the LSFS path, the directory name and filename of both files, and the connection conditions. The Default Parameters are the destination directory for file 2 and the connection condition, while the LSFS path, directory names, and filenames are Positional Parameters. If the destination directory for file 2 does not exist, the two files will be placed in the same directory. If the destination directory exists, the files will be placed in their respective directories. If the join condition is set to new, the syscall will preserve the original files, concatenate their text contents, and create a new file in the destination directory of file 1. The new name of file will be a combination of the two target filenames. If the join condition is not new, the contents of file 2 will be appended directly to the contents of file 1, and file 2 will then be deleted.
Appendix B The Implementation Details of LSFS API
Retrieve-Summary API.
Change-Summary API.
Rollback API.
This API is designed to achieve version restoration by utilizing the version recorder from the change-summary API, making version rollbacks more manageable. In this API, the overwrite() and create_or_get_file() syscalls are employed. Traditional file systems like ZFS, BtrFS, and NTFS offer rollback capabilities, but they primarily rely on snapshots, where the system captures the state of files at a specific point in time and restores them to that version. Our rollback API in the LSFS introduces two rollback methods for greater flexibility and ease of use. The first is time-based rollback, where the LLMs parses the rollback time from the input of user and reverts the file to the corresponding version. The second is version-based rollback, allowing users to specify in the prompt how many versions backward they wish to revert. This dual approach makes it easier for users to rollback to the target version of files.
Link API.
This API is designed to enable the creation of shareable file links. In traditional file systems, links can only be generated for local access, limiting collaboration. However, with the link API of LSFS, shareable links can be generated for broader accessibility. Specifically, the API leverages cloud database, e.g., Google Drive, to upload files and generate the shareable link. Additionally, validity period can be passed as an argument for this API, once the period expires, the link API automatically revokes the link and terminates access for secure and time-bound file sharing.
Appendix C The Instruction Examples of executing API
This section introduces some instruction examples of executing LSFS APIs, which can be seen at Table 5.
Table 5: Some examples of instruction of API in Section 4.3. For every API, we provide different instructions in different task condition. The instruction of retrieve-summary API is in Appendix D.
API Type Instruction
w/o directory: Modify /xxxx/xxxx.txt to contain change-file.
w/o period of validity: Generate a link for system-architecture.
Appendix D Task details of keyword-based and semantic retrieval.
In this section, Table 6 and Table 7 presents instruction with or without LSFS using LLM in semantic-based retrieval, and keyword-based retrieval (single-condition and multi-condition), respectively.
@@ Appendix C The Instruction Examples of executing API
Task Task Example Method Instruction
Keyword-based Retrieval(Single-Condition) Find papers in the computer-vision category authored by Emily Zhang. LLM w/o LSFS At current step, you need to judge if the input paper satisfy [retrieve condition]. If yes, you should summarize the paper, if no you do not need to output anything. The paper is [file content].
Keyword-based Retrieval(Multi-Condition) Find papers from either Cambridge University or Columbia University. LLM w/o LSFS At current step, you need to judge if the input paper satisfy [retrieve condition]. If yes, you should summarize the paper, if no you do not need to output anything. The paper is [file content]
Appendix E Baselines in keyword-based file retrieval and file sharing
The baselines to compare with LSFS in keyword-based file retrieval are as below:
  • TFS search window: We use the default search window in the file of computer folder to retrieve files (i.e., Spotlight in MacOS) which supports retrieving by keywords in the file content.
The baselines to compare with LSFS in file sharing are as below:
  • GPT-4o-mini: We employ GPT-4o-mini as the LLM backbone to generate code for creating links of the target file, followed by using the Python compiler to verify the validity of code.
Appendix F A Case study of Semantic File Retrieval
Figure 7: A case study of comparing semantic file retrieval between LLM-leveraged methods without using LSFS and using LSFS.
[…]
Appendix G Additional Analysis of File Sharing
In this section, we analyze the failure cases of file sharing in baselines and use some pseudo-code examples to show the file sharing process in baselines and LSFS, respectively. For all the methods, we set the following input: You are good at writing code, please write code to generate shared links for the file ’path’ as the system prompt for backbone LLMs.
G.1 The code cannot generate link
In the experiment, the code generated by LLMs may not produce the correct link or link address. For instance, even after successfully installing the required file packages, the following code block demonstrates that the generated link does not direct to the intended target file. The pseudo-code is in Algorithm 1.
Algorithm 1 Pseudo-code of G.1.
@@ G.1 The code cannot generate link
G.2 The code can only generate local link
@@
Algorithm 2 Pseudo-code of G.2.
@@
G.3 The code can generate shareable link
@@
Appendix H Further Discussions
@@
The core functionality of LSFS lies in its keyword and semantic retrieval capabilities. In future work, these retrieval methods can be further optimized by integrating more advanced and precise algorithms, enhancing the overall accuracy and effectiveness of retrieval performance of LSFS.